### v2

# Gradient-flow adaptive importance sampling for Bayesian leave one out cross-validation with application to sigmoidal classification models

Abstract
We introduce gradient-flow-guided adaptive importance sampling (IS) transformations for stabilizing Monte-Carlo approximations of leave-one-out (LOO) cross-validated predictions for Bayesian models. After defining two variational problems, we derive corresponding simple nonlinear transformations that utilize gradient information to shift a model’s pre-trained full-data posterior closer to the target LOO posterior predictive distributions. In doing so, the transformations stabilize importance weights. The resulting Monte Carlo integrals depend on Jacobian determinants with respect to the model Hessian. We derive closed-form exact formulae for these Jacobian determinants in the cases of logistic regression and shallow ReLU-activated artificial neural networks, and provide a simple approximation that sidesteps the need to compute full Hessian matrices and their spectra. We test the methodology on an $n\ll p$ dataset that is known to produce unstable LOO IS weights.
1 Introduction
In Bayesian inference, multiple models are often fit to data and then a selection procedure is applied to decide which model is best. A commonly used criterion for selection is how well a model accurately predicts out-of-sample data, which automatically optimizes fit while avoiding over-fitting. Prediction accuracy is most naturally estimated using cross-validation (CV). A model is trained with some of the data and tested against the unused data. However, estimates of out-of-sample model metrics using train-test splitting are noisy [Dietterich, 1998, Kohavi, 1995] unless computationally expensive k-fold cross-validation (i.e. fitting the model multiple times on cross-over the entire dataset) is employed [Rodriguez et al., 2010, Wong and Yeh, 2020].
Although $N$-fold, also known as leave one out (LOO), CV is the most expensive, there exist computationally efficient LOO techniques that completely avoid refitting. For example, the Akaike Information criteria (AIC) and Bayesian variants [Stone, 1977, Watanabe, 2010, Gelman et al., 2014, Watanabe, 2013] are asymptotic approximations of LOO CV. For Bayesian models, a more precise way to perform LOO CV is to use importance sampling [Vehtari et al., 2017, Piironen and Vehtari, 2017a], which works by using the full data posterior measure as a proposal distribution for each data point’s LOO posterior measure. These methods preclude refitting the model by approximating each LOO posterior measure using full data posterior measure. However, in cases where the LOO measure and full measure are very different, importance sampling can fail [Piironen and Vehtari, 2017a]. To ameliorate this possibility, we introduce an adaptive importance sampling method for LOO CV based on using transformations that bring the proposal distribution closer to LOO posteriors. We derive these transformations by defining gradient flows that minimize given statistical objective. While the transformations are model-dependent, the method is made completely general when using autograd for computing model gradients – for computational efficiency we derive the transformations exactly for a large class of classification models.
2 Preliminaries
Notation: We denote vectors (assumed to be column vectors unless otherwise stated) using bold-faced lowercase symbols, and matrices using bold-faced uppercase symbols. Given a matrix $\mathbf{W}=(w_{ij})$, the $i$-th row is denoted $\mathbf{w}_{i,}$ and $j$-th column is denoted $\mathbf{w}_{,j}$.
We refer to the entire set of observed training data as $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$. As shorthand, we denote the set of training data where the i-th observation is left out as $\mathcal{D}^{(-i)}=\mathcal{D}\setminus\{(\mathbf{x}_{i},y_{i})\}$. Expectations with respect to the posterior distribution of $\boldsymbol{\theta}$ are denoted $\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}}$, and with respect to the posterior distribution of $\boldsymbol{\theta}$ if observation $i$ is left out are denoted $\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}$.
For a transformation $T:\Omega\to\Omega,$ where $\Omega\subset\mathbb{R}^{p}$, we denote its Jacobian matrix $\mathbf{J}_{T}=\nabla T=\left(\partial_{\alpha}T^{\beta}\right)_{\alpha,\beta}$ and the determinant of the Jacobian matrix $\mathcal{J}_{T}=|\mathbf{J}_{T}|.$ The gradient operator $\nabla$ operating on a function $\mu:\mathbb{R}^{p}\to\mathbb{R}$ is assumed to yield a column vector, the Hessian matrix for a function $\mu$ is denoted $\nabla\nabla\mu,$ and the Laplacian of $\mu$ is denoted $\nabla^{2}\mu$
The operator ||$\cdot$|| ⋅ | refers to determinants when the argument is a matrix, the 2-norm when the argument is a vector, and the absolute value when the argument is a scalar.
Importance sampling-based approximate leave one out cross validation (IS-LOO): Suppose that one has pre-trained a Bayesian model such that one is able to sample its posterior parameters
Figure 1: Relationships between probability densities. One wants to sample from $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)}),$ the LOO distribution for observation $i$, by sampling from the full-data posterior $\pi(\boldsymbol{\theta}|\mathcal{D}).$ The transformation $T_{i}$ on the full-data posterior brings the sampling distribution closer to the target LOO distribution.
$\boldsymbol{\theta}_{s}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D}).$ Our objective is to use knowledge of this full-data posterior distribution to estimate how the model would behave if any single point is left out at training. One can relate the full-data model to the model with observation $i$ left out using the Bayesian update equation
$\pi(\boldsymbol{\theta}|\mathcal{D})=\frac{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})}{\int\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})\mathrm{d}\boldsymbol{\theta}},$ (1)
which is a Fredholm integral equation of thge with respect second-kind to $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$. This integral equation is in-practice difficult to solve due to the typically-high dimensionality of $\boldsymbol{\theta}.$
Rather than directly inverting Eq. 1, our starting point is the observation that Eq. 1 implies
$\frac{\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})}{\pi(\boldsymbol{\theta}|\mathcal{D})}=\frac{\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}[\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})]}{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})}\equiv\nu_{i}(\boldsymbol{\theta}),$ (2)
which provides the ratio of densities between a distribution we know (the full-data posterior $\pi(\boldsymbol{\theta}|\mathcal{D})$) and a distribution whose statistics we would like to compute (the point-wise LOO posterior $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$). To use the former to compute statistics of the latter we turn to Monte-Carlo [Barbu and Zhu, 2020, Robert and Casella, 2013] - the use of statistical sampling to compute a desired quantity (typically an integral). Importance Sampling (IS) is a Monte-Carlo method where one computes expectations with respect to a target distribution by taking a weighted average of samples with respect to a given proposal distribution. For an integrable function $f$,
$\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]=\int f(\boldsymbol{\theta})\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})\mathrm{d}\boldsymbol{\theta}$
$=\int f(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})}{\pi(\boldsymbol{\theta}|\mathcal{D})}\pi(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}=\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}}\left[f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\right].$ (3)
We approximate Eq. 3 by sampling over $\boldsymbol{\theta}_{k}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D})$, and computing the Monte-Carlo integral
$\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\approx\sum_{k=1}^{s}\nu_{ik}f(\boldsymbol{\theta}_{k})$ (4)
where the coefficients $\nu_{ik}$ are known as the self-normalized importance sampling weights
$\nu_{ik}=\frac{\nu_{i}(\boldsymbol{\theta}_{k})}{\sum_{j=1}^{s}\nu_{i}(\boldsymbol{\theta}_{j})}=\frac{(\ell(\boldsymbol{\theta}_{k}|y_{i},\mathbf{x}_{i}))^{-1}}{\sum_{k=1}^{s}(\ell(\boldsymbol{\theta}_{k}|y_{i},\mathbf{x}_{i}))^{-1}},$ (5)
so that the undetermined constant $\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})\right]$ cancels out. Eqs. 45 define a well-known [Gelfand et al., 1992] Monte-Carlo estimator for LOO.
LOO-based metrics: A Bayesian analogue to the Aikaike Information Criterion (AIC) can be computed using the LOO information criterion (LOO-IC):
LOO-IC $=-2\sum_{i=1}^{n}\log\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left(\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})\right)$
$\approx-2\sum_{i=1}^{n}\log\sum_{k=1}^{s}{\nu}_{ik}\ell(\boldsymbol{\theta}_{k}|y_{i},\mathbf{x}_{i}).$ (6)
Alternatively, for classification problems, one often desires other metrics such as an estimate of the out-of-sample area under the receiver operator curve or precision-recall curve. To compute these quantities, one simply propagates LOO estimates of the outcome probabilities
$\widehat{\mathscr{p}}_{\textrm{loo},i}=\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[\mathscr{p}_{i}(\boldsymbol{\theta})\right]\approx\sum_{k=1}^{s}{\nu}_{ik}\mathscr{p}(\boldsymbol{\theta}_{k},\mathbf{x}_{i}),$ (7)
into the relevant formulae. For instance, one may use the probabilistic interpretation of the AUROC [Wang and Guo, 2020] to motivate the direct estimator of the LOO AUROC,
[…]
where $\mathcal{D}^{0}$ is the set of all negative observations and $\mathcal{D}^{1}$ is the set of all positive observations.
Weight stabilization: Often it is the case that using the computed posterior $\pi(\boldsymbol{\theta}|\mathcal{D})$ as the proposal distribution for importance sampling has slow convergence properties – the $1/\ell$ importance weights are known to have large or unbounded variance [Peruggia, 1997], making the importance sampler estimate for LOO noisy.
Two practical model agnostic methods for controlling the tail of the importance weights are through weight truncation [Ionides, 2008] and Pareto smoothing  [Vehtari et al., 2024, 2017]. Pareto smoothing replaces the largest $M$ weights with their corresponding rank-values from a fitted generalized Pareto-distribution [Zhang and Stephens, 2009]. Pareto smoothed importance sample (PSIS)-based LOO implementations are widely available in software packages such as Stan and ArviZ. However, PSIS-LOO is insufficient in problems where the Pareto distribution does not well-fit the tail distribution of importance weights – where the estimated Pareto shape parameter $\hat{k}$ exceeds $0.7$ – this parameter is a useful criterion for accessing the validity of a given importance sampler. In such a case, it would be desirable to perform an additional model-specific controlled transformation on the proposal distribution that will induce more efficient computations.
2.1 Adaptive importance sampling
We apply the concept of adaptive importance sampling [Bugallo et al., 2017, Cornuet et al., 2011, Elvira and Martino, 2022] to transform the posterior distribution to be closer to the LOO distribution $\pi(\boldsymbol{\theta}|\mathcal{D}^{(-i)})$ (relationships between the different distributions is depicted in Fig. 1).
Consider the bijection $T_{i}:\mathbb{R}^{p}\to\mathbb{R}^{p}$, defined for observation $i,$ and let $\boldsymbol{\phi}\equiv T_{i}(\boldsymbol{\theta}).$ By change of variables, $\pi_{\boldsymbol{\phi}}(\boldsymbol{\phi}|\ldots)=\pi\left(T_{i}^{-1}(\boldsymbol{\phi})|\ldots\right)\mathcal{J}_{i}^{-1}(\phi),$ where we denote $\mathbf{J}_{T}=\nabla T$, $\mathcal{J}_{T_{i}}^{-1}(\boldsymbol{\phi})=\left|\mathbf{J}_{T_{i}}^{-1}(\boldsymbol{\phi})\right|,$ and $\mathcal{J}_{T_{i}}(\boldsymbol{\theta})=\left|\mathbf{J}_{T_{i}}(\boldsymbol{\theta})\right|={1}/{\mathcal{J}_{T_{i}}^{-1}(\boldsymbol{\phi})},$ the determinants of the Jacobians of the inverse and forward transformations respectively. One can rewrite the expectation in Eq. 3 in terms of an integral over $\pi_{\boldsymbol{\phi}}$,
${\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\pi(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}}$
$=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D})}{\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})}\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}$
$=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D})\mathcal{J}_{T_{i}}(T_{i}^{-1}(\boldsymbol{\theta}))}{\pi(T_{i}^{-1}(\boldsymbol{\theta})|\mathcal{D})}\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}.$ (9)
One can then define a Monte-Carlo approximation of Eq. 9 using importance sampling, by sampling $\boldsymbol{\theta}_{k}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D})$ so that $\boldsymbol{\phi}_{k}=T_{i}(\boldsymbol{\theta}_{k})\overset{\mathrm{iid}}{\sim}\pi_{\boldsymbol{\phi}}(\boldsymbol{\phi}|\mathcal{D}):$
$\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]$ $\approx\sum_{k=1}^{s}\frac{\eta_{ik}}{\sum_{j=1}^{s}\eta_{ij}}f(\boldsymbol{\phi}_{k})$ (10)
$\eta_{ik}$ $=\frac{\mathcal{J}_{T_{i}}(\boldsymbol{\theta}_{k})}{\ell(\boldsymbol{\phi}_{k}|\mathbf{x}_{i},y_{i})}\frac{\pi(\boldsymbol{\phi}_{k}|\mathcal{D})}{\pi(\boldsymbol{\theta}_{k}|\mathcal{D})}.$ (11)
By Bayes rule (Eq. 1), the posterior likelihood ratio in Eqs. 911 has the exact expression
$\frac{\pi(\boldsymbol{\phi}|\mathcal{D})}{\pi(\boldsymbol{\theta}|\mathcal{D})}=\prod_{i}\frac{\ell(\boldsymbol{\phi}|\mathbf{x}_{i},y_{i})}{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})}.$ (12)
Computing this expression requires iterating over the entire dataset. For large datasets, one can turn to variational approximations.
Using variational posteriors: For computational expediency, variational methods are often used in place of MCMC for Bayesian inference, obtaining a variational approximation $\hat{\pi}(\boldsymbol{\theta}|\mathcal{D})$ to the true posterior, where $\hat{\pi}$ lies within a given family of transformed probability distributions. In problems where one expects a substantial discrepancy between the true posterior and $\hat{\pi},$ one may correct for this discrepancy by noting that
${\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\pi(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}}$
$=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D})}{\hat{\pi}_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})}\hat{\pi}_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}$
$=\int f(\boldsymbol{\theta})\nu_{i}(\boldsymbol{\theta})\frac{\pi(\boldsymbol{\theta}|\mathcal{D})\mathcal{J}_{T_{i}}(T_{i}^{-1}(\boldsymbol{\theta}))}{\hat{\pi}(T_{i}^{-1}(\boldsymbol{\theta})|\mathcal{D})}\hat{\pi}_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\mathrm{d}\boldsymbol{\theta}$ (13)
and using the self-normalized importance sampler
$\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]\approx\sum_{k=1}^{s}\frac{\chi_{ik}}{\sum_{j=1}^{s}\chi_{ij}}f(\boldsymbol{\phi}_{k})$ (14)
$\qquad\chi_{ik}=\frac{\mathcal{J}_{i}(\boldsymbol{\theta}_{k})}{\hat{\pi}(\boldsymbol{\theta}_{k}|\mathcal{D})}\pi(\boldsymbol{\phi}_{k})\prod_{j\neq i}\ell(\boldsymbol{\phi}_{k}|\mathbf{x}_{j},y_{j}),$ (15)
where $\pi(\boldsymbol{\phi}_{k})$ is the prior density at $\boldsymbol{\phi}_{k}$, canceling out the two unknown constants corresponding to $\pi(\boldsymbol{\phi}_{k}|\mathcal{D})$ and $\nu_{i}.$
3 Methods
Our focus is on classification models where a vector of covariates $\mathbf{x}\in\mathbb{R}^{p}$ is used to estimate the probability of an outcome labeled by $y\in\{0,1\}$ with likelihood function $\ell$:
$y_{i}|\boldsymbol{\theta},\mathbf{x}_{i}$ $\ {\sim}\ \textrm{Bernoulli}\left(\mathscr{p}_{i}(\boldsymbol{\theta})\right)$
$\ell(\boldsymbol{\theta}|y_{i},\mathbf{x}_{i})$ $=\mathscr{p}_{i}(\boldsymbol{\theta})^{y_{i}}(1-\mathscr{p}_{i}(\boldsymbol{\theta}))^{1-y_{i}},$ (16)
and where $\mathscr{p}_{i}(\boldsymbol{\theta})\equiv\mathscr{p}(\boldsymbol{\theta},\mathbf{x}_{i})$ is the predicted outcome probability for observation $i$, and computes the statistics of the parameter vector $\boldsymbol{\theta}$ by applying Bayes rule
$\pi(\boldsymbol{\theta}|\mathcal{D})\propto\pi(\boldsymbol{\theta})\prod_{i=1}^{n}\ell(\boldsymbol{\theta}|y_{i},\mathbf{x}_{i}),$ (17)
where $\pi(\boldsymbol{\theta})$ is the joint prior distribution over the parameter vector $\boldsymbol{\theta}$, corresponding to the likelihood function $\ell(\boldsymbol{\theta})$. There are usually many plausible choices for $\ell,$ even within the same family of models (for example different sets of features in logistic regression) – one needs a method to evaluate different models to inform decisions such as model selection.
Eq. 11 is valid for an arbitrary bijection $T_{i}.$ The objective of using transformations is to shift the proposal distribution closer to the targeted LOO distribution for each observation – to partially invert the Bayes rule. To this end, we motivate two different transformations of the form
$T_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}+hQ_{i}(\boldsymbol{\theta}),$ (18)
for some small step-size $h$ and function $Q_{i}.$
3.1 Gradient flow transformations
KL divergence descent: We consider choosing $T_{i}$ to minimize the KL divergence $\mathrm{D}_{\mathrm{KL}}\left(\pi({\boldsymbol{\theta}|\mathcal{D}^{(-i)}})\|\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\right)$, which is equivalent to minimizing the cross-entropy with respect to the mapping $T_{i}$,
$H\left(\pi({\boldsymbol{\theta}|\mathcal{D}^{(-i)}}),\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\right)=$
$\qquad-\int\nu_{i}(\boldsymbol{\phi})\pi(\boldsymbol{\phi}|\mathcal{D})\log\frac{\pi(T_{i}^{-1}(\boldsymbol{\phi})|\mathcal{D})}{\mathcal{J}_{T_{i}}(T_{i}^{-1}(\boldsymbol{\phi}))}\mathrm{d}\boldsymbol{\phi}.$ (19)
The Euler-Lagrange equation for minimizing Eq. 19 (derived in Supplemental Materials S.1.1), is implicit in $T_{i}.$ While it admits no closed form solution, one may note that $T_{i}$ is a $t\to\infty$ stable fixed point of the KL-descending gradient flow
$\frac{\partial T_{i}(\boldsymbol{\theta},t)}{\partial t}=-\frac{\delta H\left(\pi({\boldsymbol{\theta}|\mathcal{D}^{(-i)}}),\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\right)}{\delta T_{i}}$ (20)
and use this fact to refine, using the method of lines, an initial guess of $T_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}$ with forward Euler discretization of step-size $h[\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}[\ell(\boldsymbol{\theta},\mathbf{x}_{i},y_{i})]]^{-1}$, for $0<h\ll 1$, to arrive at the transformation
$T^{\textrm{KL}}_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}-h\frac{\delta H\left(\pi({\boldsymbol{\theta}|\mathcal{D}^{(-i)}}),\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\right)}{\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}[\ell(\boldsymbol{\theta},\mathbf{x}_{i},y_{i})]\delta T_{i}}\Bigg{|}_{T(\boldsymbol{\theta})=\boldsymbol{\theta}}$
$=\boldsymbol{\theta}+h\underbrace{\pi({\boldsymbol{\theta}}|\mathcal{D})\nabla\left(\frac{1}{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})}\right)}_{Q_{i}^{\textrm{KL}}}.$ (21)
Variance descent: In importance sampling, the variance of the estimator is conditional on the target function for expectation. Since we are interested in computing the LOO predictive probability for each observation $i$, it is natural to consider minimizing the variance of the transformed importance sampler for the function $\mathscr{p}_{i}(\boldsymbol{\theta})=\mathscr{p}(\boldsymbol{\theta},\mathbf{x}_{i}).$ However, this objective yields a transformation that is only useful for observations where $y_{i}=0$ (see Supplemental MaterialsS.1.2). Instead, we seek to minimize the variance with respect to estimating the complement probability $\mathscr{p}_{i}(\boldsymbol{\theta})^{1-y_{i}}(1-\mathscr{p}_{i}(\boldsymbol{\theta}))^{y_{i}}.$
Starting from the associated variational problem (Appendix S.1.2), and applying the same rationale that went into developing the KL-descending transformation, one arrives at the single-step variance-reducing transformation,
$T^{\textrm{Var}}_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}+hQ_{i}^{\textrm{Var}}(\boldsymbol{\theta})$
$Q_{i}^{\textrm{Var}}(\boldsymbol{\theta})$ $=\pi(\boldsymbol{\theta}|\mathcal{D})\left[\frac{1-\mathscr{p}_{i}(\boldsymbol{\theta})}{\mathscr{p}_{i}(\boldsymbol{\theta})}\right]^{2y_{i}-1}$
×$\qquad\times\nabla\left(\left[\frac{1-\mathscr{p}_{i}(\boldsymbol{\theta})}{\mathscr{p}_{i}(\boldsymbol{\theta})}\right]^{2y_{i}-1}\right).$ (22)
Resolving the posterior density: Both the KL (Eq. 21) and variance (Eq. 22) descent transformations take steps proportional to the posterior density $\pi(\boldsymbol{\theta}|\mathcal{D}).$ If a variational approximation for $\pi({\boldsymbol{\theta}}|\mathcal{D})$ is available, using it in Eqs. 21 and 22 as a stand-in for the posterior density helps simplify the computation of the transformations and their Jacobians, particularly when using mean-field or low-order Automatic Differentiation Variational Inference (ADVI) [Kucukelbir et al., 2017, Blei et al., 2017].
In the absence of variational approximation, one may evaluate the posterior densities exactly using the Bayes rule, absorbing the unknown normalization constant $\mathcal{Z}$ into the step size $h.$ Then, the KL and variance descent transformations are
$T^{\textrm{KL}}_{i}(\boldsymbol{\theta})$
$=\boldsymbol{\theta}+h\pi(\boldsymbol{\theta})\left[\prod_{j}\ell(\boldsymbol{\theta}|\mathbf{x}_{j},y_{j})\right]\nabla\left(\frac{1}{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})}\right),$ (23)
and $T^{\textrm{Var}}_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}+hQ_{i}^{\textrm{Var}}(\boldsymbol{\theta}),$ where
$Q_{i}^{\textrm{Var}}(\boldsymbol{\theta})$ $=\pi(\boldsymbol{\theta})\left[\prod_{j}\ell(\boldsymbol{\theta}|\mathbf{x}_{j},y_{j})\right]\left[\frac{1-\mathscr{p}_{i}(\boldsymbol{\theta})}{\mathscr{p}_{i}(\boldsymbol{\theta})}\right]^{2y_{i}-1}$
×$\qquad\times\nabla\left(\left[\frac{1-\mathscr{p}_{i}(\boldsymbol{\theta})}{\mathscr{p}_{i}(\boldsymbol{\theta})}\right]^{2y_{i}-1}\right).$ (24)
The obvious downside of using these exact transformations is the need to iterate over the entire dataset in order to evaluate the posterior density, which must be done for each parameter sample, for each data point.
For evaluating the Jacobian determinants, one appeals to Bayes rule to find that
$\nabla\log[\mathcal{Z}\pi(\boldsymbol{\theta}|\mathcal{D})]=\nabla\log\pi(\boldsymbol{\theta})+\sum_{i}\nabla\log\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i}),$ (25)
where $\mathcal{Z}$ is is absorbed into $h$.
Step size selection The KL-divergence and variance descent transformations correspond to a forward Euler solver on the respective gradient flow equations. According to linear stability analysis, Euler’s method has the conditional stability criteria $h<{2}/{\max_{k}|\mathrm{Re}(\lambda_{k})|}$ where $\lambda_{k}$ are the eigenvalues of the Jacobian of the system (Jacobians of the functions $Q_{i}$). In each case the structure of the Jacobian admits cheap approximations of $\lambda_{k}$. However, for nonlinear systems, this criterion is not sufficient for achieving stability.
Instead, we use a modified rule to determine the step size. For all parameter samples at each individual observation $i,$ we use
$h_{i}=\rho\min_{s,\alpha}\left\{\left|{\sqrt{\Sigma_{\alpha,\alpha}}}/{Q_{i}(\boldsymbol{\theta}_{s})_{\alpha}}\right|\right\}$ (26)
where $\rho>0$ and $\sqrt{\Sigma_{\alpha,\alpha}}$ is the marginal posterior standard deviation of the $\alpha$-th component of $\boldsymbol{\theta}.$ This rule ensures that the transformation takes a step of at most $\rho$ posterior standard deviations in any parameter component. The objective of adaptation is to find any transformation that results in importance weights where the Pareto tail shape is sub-threshold. To this end, one can compute the transformations for a range of $\rho$ values in parallel using vectorized computations, saving computation at the cost of memory utilization.
Jacobian determinant approximation: For either single-step transformations, one may approximate $\left|\mathbf{J}_{T_{i}}\right|$ by noting that
$\mathcal{J}_{T_{i}}(\boldsymbol{\theta})=\left|1+h\nabla\cdot Q_{i}(\boldsymbol{\theta})\right|+\mathcal{O}(h^{2})$ (27)
and truncating to $\mathcal{O}(h),$ sidestepping the computation of Hessian matrices and their spectra. Note that any higher order terms in this expansion require characterization of the spectra of $\nabla Q_{i},$ for each observation $i$, and for each sampled parameter $\boldsymbol{\theta}_{k}$. For large problems, computing the Jacobian matrix and its spectra many times can become computationally problematic.
Overview: We have presented two gradient-flow descending transformations, each aimed at stabilizing a LOO importance sampler by bringing the proposal distribution closer to the LOO target in a different sense. The KL divergence and variance descent transformations correspond to the first-step of a forward Euler discretized solver for the corresponding gradient flow equations. The log-likelihood descent transformation repels the posterior parameter distribution away from the targeted LOO observation, undoing part of its contribution to the full data posterior. While each transformation uses gradient information, their Jacobians are simple to approximate, requiring no computation of full Hessian matrices.
Generally, one will find that many observations are amenable to direct importance sampling with $1/\ell$ weights (Eq. 3) in combination with Pareto smoothing (tail weight distribution shape parameter $\hat{k}<0.7$). One needs only transform the sampling distribution when the estimated shape parameter exceeds this threshold. For a given posterior sample of model parameters $\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{s}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D})$, one undergoes the procedure for each given observation $i$:
procedure $\mathbf{AdaptiveIS}$(observation $i$)
     Compute weights $\nu_{ik}$ (Eq. 5) and their tail shape $\hat{k}$
     if $\hat{k}\leq 0.7$ then
         Done
     for $T_{i}$ in transformations do
         Apply $T_{i}$ to each $\boldsymbol{\theta}_{k}$
         Compute weights $\eta_{ik}$ (Eq. 11)
         Compute $\hat{k}$
         if $\hat{k}\leq 0.7$ then
              Done
It is important to note that if any transformation takes $\hat{k}$ for a given observation under the threshold then adaptation is successful.
4 Examples
In this manuscript, we consider the broad widely-used class of models that have a sigmoidal parameterization
$\mathscr{p}_{i}(\boldsymbol{\theta})=p(\boldsymbol{\theta},\mathbf{x}_{i})=\sigma(\mu_{i}(\boldsymbol{\theta}))$ (28)
where $\sigma(\mu)=1/(1+e^{-\mu})$ is the sigmoid function and we denote $\mu_{i}(\boldsymbol{\theta})\equiv\mu(\boldsymbol{\theta},\mathbf{x}_{i})$ for some mean function $\mu$.
For these models, the transformations take the form
$Q_{i}^{\textrm{KL}}(\boldsymbol{\theta})$ $=(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\nabla\mu_{i}$ (29)
$Q_{i}^{\textrm{Var}}(\boldsymbol{\theta})$ $=(-1)^{y_{i}}\pi(\boldsymbol{\theta}|\mathcal{D})e^{2\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\nabla\mu_{i},$ (30)
and their Jacobians take the form
$\mathbf{J}_{T_{i}^{\textrm{KL}}}(\boldsymbol{\theta})=\mathbf{I}+h(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\Big{\{}\nabla\nabla\mu_{i}$
$\quad+\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+(1-2y_{i})\nabla\mu_{i}\right](\nabla\mu_{i})^{\intercal}\Big{\}}$ (32)
$\mathbf{J}_{T_{i}^{\textrm{Var}}}(\boldsymbol{\theta})=\mathbf{I}+h(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{2\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\Big{\{}\nabla\nabla\mu_{i}$
$\quad+\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+2(1-2y_{i})\nabla\mu_{i}\right](\nabla\mu_{i})^{\intercal}\Big{\}},$ (33)
where
$\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})=\nabla\log\pi(\boldsymbol{\theta})$
$+\sum_{j}\left(y_{j}(1-\sigma(\mu_{j}))-(1-y_{j})\sigma(\mu_{j})\right)\nabla\mu_{j}(\boldsymbol{\theta}),$ (34)
and $\pi(\boldsymbol{\theta})$ is the prior. Here we will consider two popular sub-families of sigmoidal models.
Logistic Regression (LR): LR is a sigmoidal model where $\mu_{i}(\boldsymbol{\theta})=\mathbf{x}_{i}^{\intercal}{\boldsymbol{\beta}}$, So, $\nabla_{\boldsymbol{\beta}}\mu_{i}=\mathbf{x}_{i}$, and $\nabla\nabla\mu=\mathbf{0}$. Because the Hessian of $\mu$ vanishes, the Jacobian of the function $Q_{i}$ for each of the functions is a rank-one matrix and has only a single non-zero eigenvalue. LR admits exact Jacobian determinants for each of the transformations:
$\mathcal{J}_{T^{\textrm{KL}}_{i}}(\boldsymbol{\theta})=\Big{|}1+h(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}$
×$\quad\quad\quad\quad\quad\times\mathbf{x}_{i}^{\intercal}\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+(1-2y_{i})\mathbf{x}_{i}\right]\Big{|},$ (35)
$\mathcal{J}_{T_{i}^{\textrm{Var}}}(\boldsymbol{\theta})=\Big{|}1+h(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{2\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}$
×$\quad\quad\quad\quad\quad\times\mathbf{x}_{i}^{\intercal}\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+2(1-2y_{i})\mathbf{x}_{i}\right]\Big{|}.$ (36)
Bayesian (ReLU) Neural Networks: Bayesian ReLU-nets [Lee, 2000, Ghosh and Doshi-Velez, 2017, Choi et al., 2018, Kristiadi et al., 2020, Bhadra et al., 2019] are piecewise linear [Sudjianto et al., 2021, Wang, 2022, Montúfar et al., 2014, Sudjianto et al., 2020] extensions to regression models. Being locally linear, these models have block-sparse Hessians and are also amenable to some limited degree of interpretability [Sudjianto et al., 2020, Chang et al., 2023]. One may write an $L$-layer ReLU Bayesian neural network recursively
$y_{i}|\mu_{i}$ $\sim\textrm{Bernoulli}(\sigma(\mu_{i}))$
$\mu_{i}|\mathbf{W}_{L},b_{L},\mathbf{z}^{(i)}_{L-1}$ $=\mu(\mathbf{x}_{i})=\mathbf{W}_{L}a(\mathbf{z}^{(i)}_{L-1})+b_{L}$
$\mathbf{z}_{k}|\mathbf{z}^{(i)}_{k-1},\mathbf{b}_{k},\mathbf{W}_{k}$ $=\mathbf{W}_{k}a(\mathbf{z}^{(i)}_{k-1})+\mathbf{b}_{k}$
$\mathbf{z}^{(i)}_{1}|\mathbf{W}_{1},\mathbf{x}_{i}$ $=\mathbf{W}_{1}\mathbf{x}_{i},$ (37)
where $a$ is the ReLU activation function. The derivative of this function is the unit step function. We assume that the output function is sigmoid, noting that the softmax function also transforms into a sigmoid under a change of variables. Within the parameterization of Eq. 37 we absorbed the initial first-layer bias into the transformation $\mathbf{W}_{1},$ by assuming that $\mathbf{x}$ has a unit constant component, as is the convention in regression.
The Hessian matrix of $\mu$, while non-zero, is sparse because all of the following identities hold:
$\nabla_{\mathbf{b}_{k}}\nabla_{\mathbf{b}_{j}}\mu$ $=\mathbf{0}\qquad\forall j,k$ (38)
$\nabla_{\mathbf{W}_{k}}\nabla_{\mathbf{W}_{k}}\mu$ $=\mathbf{0}\qquad\forall k$ (39)
$\nabla_{\mathbf{W}_{k}}\nabla_{\mathbf{b}_{j}}\mu$ $=\mathbf{0}\qquad\forall j\geq k.$ (40)
For this reason, the Jacobian determinant approximation of Eq. 27 can ignore the model Hessian entirely. However, in the case of one hidden layer we exploit the Hessian’s structure to provide explicit exact expressions for $\mathcal{J}_{(\cdot)}.$
Example 4.1 (One hidden layer).
These models are governed by the equations $\mu=\mathbf{W}_{2}a(\mathbf{z}_{1})+b_{2}$ and $\mathbf{z}_{1}=\mathbf{W}_{1}\mathbf{x}$, where $\mathbf{W}_{2}\in\mathbb{R}^{1\times d},$ $b_{2}\in\mathbb{R}$, $\mathbf{W}_{1}\in\mathbb{R}^{d\times p},$ $\mathbf{b}_{1}\in\mathbb{R}^{d}.$ This model has the first-order derivatives ${\partial_{({W}_{2})_{1i}}}\mu=a(z_{1})_{i}$, ${\partial_{({W}_{1})_{ij}}}\mu=(W_{2})_{1i}a^{\prime}((z_{1})_{i})x_{j}$, ${\partial_{b_{2}}}\mu=1.$ The only non-zero components of the Hessian matrix for $\mu$ are the mixed partial derivatives
$\frac{\partial^{2}\mu}{\partial{(W_{1})_{jk}}\partial({W}_{2})_{1j}}=a^{\prime}((z_{1})_{j})x_{k}.$ (41)
The Hessian matrix of $\mu$ has a particular block structure that can be exploited (see Supplemental Materials  S.2.1.1 for derivations) in order to find explicit expressions for its $2d$ non-zero eigenvalues, for $k\in\{1,2\ldots,d\}$,
$\lambda_{k}^{\pm}=\pm\left[{\sum_{j}a^{\prime}((z_{1})_{k})x_{j}^{2}}\right]^{1/2},$ (42)
and associated eigenvectors
$\mathbf{v}_{k}^{\pm}=\left(\begin{matrix}\tilde{\mathbf{u}}_{k}/\sqrt{2|\mathbf{u}_{k}|^{2}}&\pm\mathbf{e}_{k}/\sqrt{2}&0\end{matrix}\right)^{\intercal},$ (43)
where
$\tilde{\mathbf{u}}_{k}=\big{(}\overbrace{0\quad\ldots\quad 0}^{(k-1)p\textrm{ zeros}}\quad\mathbf{u}_{k}^{\intercal}\quad\overbrace{0\quad\ldots\quad 0}^{(d-k)p\textrm{ zeros}}\big{)}^{\intercal},$ (44)
and $\mathbf{u}_{k}=a^{\prime}((z_{1})_{k})\mathbf{x}.$ To compute the overall transformation Jacobians, one can then apply rank-one updates to $\nabla\nabla\mu$ – a process that is aided by projecting the model gradients into the eigenspace of the model Hessian (see S.2.1 for derivations).
5 Experiments
PDF versions of Jupyter notebooks for producing the results in the text are included in the Supplemental Results. As baselines for comparison, we also evaluated generalizations of the moment matching (MM1/MM2) affine transformation method of Paananen et al. [2021] where we used a pre-factor of $\rho$ to control the magnitude of the change in the transformation (details in  S.3), and the log-likelihood (LL) gradient descent method of Elvira et al. [2022] (derivations of this transformation for sigmoidal models are available in  S.1).
Dataset and model: For demonstration, we used a public domain ovarian cancer micro-array dataset Hernández-Lobato et al. [2010], Schummer et al. [1999]. This dataset consists of $n=54$ observations of $p=1056+1$ predictors. As an example of a $p\gg n$ problem, model-agnostic $1/\ell$ importance sampling is insufficient for computing LOO expectations. Notably, Paananen et al. [2021] used this dataset to test their moment-matching adaptive importance sampler where they successfully decreased the number of observations where $\hat{k}>0.7$ from approximately $35$ to approximately $20$ using $s=1000$ posterior samples. We reproduced their logistic regression model, using the same regularized-horseshoe [Piironen and Vehtari, 2017b, c, Carvalho et al., 2009] prior, and the same statistical inference scheme within Stan, which we interfaced to Python using the package cmdstanpy. We ran four parallel Markov Chains, with twelve thousand burn-in iterations, retaining 2000 samples per chain (more details available in  S.4.1). We then evaluated the transformation methods on resamplings of the retained MCMC samples.
Adaptation:
KL Var LL MM1 MM2 #u̇nsuccessful
$31.5\pm 2.8$
$16.2\pm 3.0$
$18.2\pm 2.6$
$12.41\pm 2.6$
$20.4\pm 3.1$
$2.9\pm 1.6$
$3.1\pm 3.1$
$0.8\pm 1.0$
$0.2\pm 0.1$
Table 1: Counts of unsuccessful adaptations (mean $\pm$ standard deviation) when using at least one of the given combination of transformations across the step sizes $\rho\in\{4^{-r}:r\in\{0,1,\ldots,10\}\}$, as seen in one hundred simulations of parameter sample size $s=1000.$ Lower is better.
[…]
We scanned different values of $\rho=4^{-r}$, for $r\in\{0,1,2,\ldots,10\}$, evaluating both transformations (KL/Var), and the comparison methods (LL/MM1/MM2) for a given value of $\rho.$ We performed this procedure $100$ times, using samples of size $s=1000.$ Recall that adaptation is successful if any of the considered transformations can reduce $\hat{k}$ to below $0.7$.
Table 1 presents statistics (mean $\pm$ standard deviation) for the number of observations where adaptation fails when using the given combination of methods. For a representative instance of the simulation procedure, Fig. 2 depicts the computed values of $\hat{k}$ obtained, organized by the index of each relevant observation within the dataset. The large black dot is the original $\hat{k}$ value and the smaller colored dots are the post-transformation KL and Var values for varying values of $\rho$. Fig. 3 shows the corresponding LOO- receiver operator characteristic (ROC) and precision recall curve (PRC) obtained by using the transformation that resulted in the best $\hat{k}$ value for each observation, predicting the LOO estimate of predictive probability, and feeding those probabilities into the relevant formulae for computing ROC and PRC. We provide these curves for both MCMC and ADVI -inferred variants of the model.
Figure 3: LOO ROC curves for ovarian cancer classification model contrasting the model fitted using MCMC and the model fitted using mean-field ADVI.
6 Discussion
In this manuscript, we introduced an adaptive importance sampler for using pre-trained full-data posteriors to approximate leave one out cross validation (LOO CV) in Bayesian classification models. The objective of importance-sampling LOO (IS LOO) is to compute LOO CV without incurring the computational cost of refitting a given model. The objective of adaptation to bring the sampling distribution (the full data posterior) closer to the target LOO posterior distributions for each data point so that IS LOO produces reliable estimates.
Our methodology is based on taking a single step according to the gradient flow corresponding to the minimization of a given objective. We introduce two such transformations: KL divergence descent and variance descent. We presented explicit formulae for these transformations for logistic regression and ReLU-activated artificial neural networks with one hidden layer – the latter by computing the exact spectral decomposition of its Hessian matrix. An analysis of the same dataset using a Bayesian ReLU-net is available in S.4.2. We described how one can easily approximate the Jacobian of the transformations for more-complicated models, including for ReLU neural networks of any size. The adaptive importance sampler is ultimately used to estimate the expected LOO prediction for each given datapoint – quantities that can be used to compute downstream model generalization metrics such as ROC/PRC curves and the area under these curves.
Contrasting and synergizing methods: Examining Tabel 1, taken individually, the KL and Var gradient flow-based transformations perform comparably to the original MM transformations evaluated in Paananen et al. [2021]. However, the generalized MM1/MM2 transformations have by-far the best performance in shifting $\hat{k}$. Yet these two MM methods, used either alone or together, usually do not completely get the job done. Each unsuccessful adaptation means that the model must be refit one additional time at high computational cost. Fortunately, using all the evaluated methods in-unison resulted in successful adaptation for all data points most of the time. The general strategy is then to loop through observations and try successive transformations for each observation until adaptation is successful.
Limitations: The main tradeoff of this method versus the model-agnostic PSIS-LOO method is that this method is model-dependent. In order to use this methodology for a given model, one needs to be able to evaluate gradients of the model with respect to parameters – and also the gradients of the corresponding prior distribution. Both the KL descent and variance descent transformations require computing the the posterior density – when a variational approximation of the posterior is not available or trustworthy this computation is costly for large datasets.
Extensions: In this manuscript, we focused on classification problems but the methodology for adapting the importance sampler is much broader. In the Supplemental Materials one may find more-general formulae for the KL and variance descending transformations. In medical and industrial contexts, one is often interested in whether an individual or unit will experience an outcome within a certain time interval. For instance, policymakers are interested in hospital readmission within 30 days post discharge [Xia et al., 2023, Chang et al., 2023] because these readmissions are possibly preventable. In these problems, one may apply survival modeling to characterize the lifetime distribution, and additionally evaluate a model according to its classification performance at a given cut-off time $T$. Our methodology can easily be used for assessing such models.
Another extension to this methodology is to take more steps along the gradient flow for a given objective. It may be feasible to learn such a transformation using neural or other expressive representations. Finally, we resolved the spectrum of the Hessian matrix for shallow ReLU models – the spectral decomposition for larger ReLU models may be useful for other analyses.
Acknowledgements
This work is supported by the Intramural Research Programs of the National Institutes of Health Clinical Center (CC) and the National Institute of Diabetes and Digestive and Kidney Diseases (NIDDK), and the US Social Security Administration.